OpenStudy (anonymous):

fan+medal. Match each shape with its perimeter or circumference. Use three point one four for pi and round to the nearest whole number where necessary. 7cm,3cm,4cm A)31cm 5cm B)34cm 12cm,5cm C)22cm 4cm D)24cm

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@miraclemilktea2

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@miraclemilktea2

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

There has to be more to this question...

OpenStudy (anonymous):

u have to match the shapes

OpenStudy (anonymous):

There are no shapes lol.

OpenStudy (anonymous):

ill past them

OpenStudy (anonymous):

Hint: The 4cm and 5cm are the circular pictures.

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

first shape is a parallelogram. perimeter of parallelogram is : P = 2(a + b) so, P = 2(7 + 4) P = 2(11) P = 22 7 cm, 3 cm, 4 cm : 22 cm (C) ================ circumference of a circle C = 2 (pi) * r C = 2(3.14)(5) C = 2(15.7) C = 31.4...rounds to 31 cm 5 cm : 31 cm (A) ================ perimeter of rectangle P = 2(l + w) P = 2(12 + 5) P = 2(17) P = 34 12 cm, 5 cm : 34 cm (B) =============== perimeter of po;ygon P = number of sides * length of a side P = 6(a) P = 6(4) P = 24 4 cm : 24 cm (D) ================= any questions ?

OpenStudy (anonymous):

thank u @texaschic101